Shift report 20030116 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Jan 16 2003 - 16:46:01 EST

  • Next message: Flemming Videbaek: "Fw: Phobos paper on comparison with e+e- multiplicity"
    ------------------------------------------------------------
    Date: Thu Jan 16 16:46:00 2003, Shift: 08:00-16:00
    Supervisor: JJG, with: 
    Log book#13 pages 68 - 72
    
    Shift Summary:
    ------------------------------------------------------------
    1. Physics run 7411 ended at around 11 am with DAQ crash. Since then no useful data due to several DAQ problems. 
    
    2. Most data since run 7410 are no good due to camac Fb3 problem =>no DC data from 7410. Problem should be fixed now by recycling the crate (switch on top of rack near TV) and the low power. But cannot really check as beam is down now. 
    
    3. Addtional problems with missing triggers due to DAQ status (must stop and restart IoServer and RunServer). 
    
    4. Run 7416 is pedestal run. 
    
    5. Discussed with Pawel and decided to do a DC HV scan: up 10,20,30 Volts and down the
    same. Right now all anodes are turned off, but demand voltages are +10 Volts, ready for beam. 
    
    6. First 16 channels of H1 slat distribution alternate in
    abnormally in yield. TRIGGER BOARD problem??? EXPERT MUST INVESTIGATE. 
    
    7. H2: TDCTOP ch 31 is dead. 
    
    8. Dana would like TOFW gate in MRS trigger
    to be 30 ns wide in the triple overlap coincidence. Need beam.
    
    9. Store dumped around 3pm.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Thu Jan 16 2003 - 16:47:00 EST